3. (10 points) Given that f(1) = 5, f(1) = 4, 9(1) = 2, and g(1) = 3, find alat (f(z)g(x)) = and d ( f(x) dx
Answer #1

Solution: Given the following values:

f(1)=5,\ f'(1)=4, \ g(1)=2,\ g'(1)=3

Calculating:

\frac{d}{dx}(f(x)g(x))|_{x=1}

Using the product rule of differentiation, we'll get:

=g(x)\frac{d(f(x))}{dx}+f(x)\frac{d(g(x))}{dx}

=g(x)\cdot f'(x)+f(x)\cdot g'(x)

Substituting x = 1, we'll get:

=g(1)\cdot f'(1)+f(1)\cdot g'(1)

Substituting the given value, we'll get:

=2\cdot 4+5\cdot 3

\Rightarrow \frac{d}{dx}(f(x)g(x))=23

Calculating:

\frac{d}{dx}(\frac{f(x)}{g(x)})|_{x=1}

Using the quotient rule of differentiation, we'll get:

=\frac{g(x)\cdot f'(x)-f(x)\cdot g'(x)}{(g(x))^2}

Substituting x = 1, we'll get:

=\frac{g(1)\cdot f'(1)-f(1)\cdot g'(1)}{(g(1))^2}

Substituting the given value, we'll get:

=\frac{2\cdot 4-5\cdot 3}{2^2}

\Rightarrow \frac{d}{dx}(\frac{f(x)}{g(x)})=-\frac{7}{4}

I hope it helps you!